A body is projected with a speed u m/s at angle β with the horizontal. The kinetic energy at the highest point is 3/4th of the initial kinetic energy. The value of β as?

Grade:11

1 Answers

Mahima Kanawat
5 years ago
Dear student... KEi=1/2mu² ; (KE)highest point =1/2mu²cos²β, Ac to given condition 1/2mu²cos²β = 3/4×1/2 mu² ; cosβ=√3/2 and hence β= 30°... WITH REGARDS, MAHIMA (ASKIITIANS FORUM EXPERT)